# On nondivergence form linear parabolic and elliptic equations with degenerate coefficients

### v1

@@ Abstract.
Corresponding author. J. Ryu was supported by a KIAS Individual Grant (MG101501) at Korea Institute for Advanced Study.
@@ 1. Introduction
Observe that equations (1.1) and (1.2) are stated without boundary conditions, as their solvability in weighted Sobolev spaces does not require such conditions. Nevertheless, it is possible to study these equations under certain boundary conditions. In such cases, additional assumptions on the forcing term $f$ or on the structure of the equations are necessary (see e.g. [40, Remark 4.7]). In this paper, however, our goal is to address the equations under minimal assumptions.
@@
where $\alpha>0$, and $H:\Omega_{T}\times\mathbb{R}^{d}\to\mathbb{R}$ is a given smooth Hamiltonian. Here, equation (1.1) corresponds to the special case $H=0$ and $\alpha=2$. We refer the reader to [14] for more information. Furthermore, the equations are derived from a variety of problems. For instance, (1.2) is obtained in the linearization of the nonlinear and degenerate Loewner–Nirenberg problem
@@
The paper is organized as follows. In Section 2, we introduce the weighted Sobolev spaces, assumptions, and our main results. In Section 3, we provide the proofs of the main results on the parabolic problem (Theorems 2.4 and 2.6). In Section 4, we prove Theorem 2.9, which concerns the elliptic problem. Lastly, we address the optimality of our main results in Section 5.
@@ 2.1. Notation and function spaces
We remark that $H_{p,\theta}^{n}$ corresponds to $H_{p,\theta+d-1}^{n}$ in [28]. Moreover, from [28, Corollary 2.3], the norm (2.1) is equivalent to
@@
To the best of our knowledge, the spaces $H_{p,\theta}^{n}$ were firstly introduced in [36, Section 2.6.3] for $p=2$ and $\theta=1$. They were generalized in a unified manner for $p\in(1,\infty)$ and $\theta\in\mathbb{R}$ in [28] to study stochastic partial differential equations (SPDEs). We refer the reader to [27, 31, 32]. See also [23, 26, 7, 39, 35], where second-order nondegenerate equations are studied.
@@
By [28, Theorem 1.19, Remark 5.5], the spaces $C_{c}^{\infty}(\mathbb{R}^{d}_{+})$, $C_{c}^{\infty}((-\infty,T)\times\mathbb{R}^{d}_{+})$, and $C_{c}^{\infty}((-\infty,T]\times\mathbb{R}^{d}_{+})$ are dense in $H_{p,\theta}^{n}$, $\mathbb{H}_{q,p,\theta,\omega}^{n}(T)$, and $\mathfrak{H}_{q,p,\theta,\omega}^{2}(T)$, respectively. For the Cauchy problems, the space $\mathring{\mathfrak{H}}_{q,p,\theta,\omega}^{2}(0,T)$ is defined by the closure of the set of functions $u\in C_{c}^{\infty}([0,T]\times\mathbb{R}^{d}_{+})$ with $u(0,\cdot)=0$, equipped with the norm
@@ 2.2. Parabolic equations
$(i)$ In [15, Theorem 2.6 $(ii)$], the corresponding divergence form equation is studied under the assumption that either
@@
$(ii)$ In Theorem 2.4 $(i)$, the lower order coefficients $(b_{i})$ and $c$ are assumed to be merely bounded measurable, and $\theta\in\mathbb{R}$ is arbitrary, although we need to choose a sufficiently large $\lambda_{0}\geq 0$. On the other hand, in the second result, we can take $\lambda=0$ by imposing ratio conditions on the lower-order coefficients and restricting the range of $\theta$.
@@ 2.3. Elliptic equations
$(ii)$ Let $n_{b},n_{c}\in\mathbb{R}$. Suppose that the quadratic equation
$z^{2}+(1+n_{b})z-n_{c}=0$ (2.11)
has two distinct real roots $\alpha<\beta$. Then there exist
@@ 3. Parabolic equations
In this section, we deal with parabolic equations and give the proofs of Theorems 2.4 and 2.6.
@@
Using (2.2), one can show that there is $R_{0}>0$ such that the (unweighted) oscillation of $x_{d}^{2}a_{ij,r}$ on $Q_{R}(t,x):=(t-R^{2},t)\times B_{R}(x)$ is less than $N\gamma_{0}$ if $x_{d}\in(1,4)$ and $R\leq R_{0}$. By [8, Theorem 6.3], there is $\Lambda_{0}=\Lambda_{0}(d,p,\nu,K_{0})\geq 0$ such that if $\Lambda=\Lambda_{0}$, then we can apply the $W_{p,\omega}^{1,2}$-estimate for the equation (3.2) to get
@@
where $L_{p,\omega}(T):=L_{p}((-\infty,T)\times\mathbb{R}^{d},\omega(t)dtdx)$. Here, we remark that the arguments in [8] still work for equations with general coefficients $a_{0}$ and $c_{0}$, and thus we can still utilize the estimate established in the paper (see also [15, Remark 3.7]). Note that for any $\gamma\in\mathbb{R}$,
@@
(ii) Let $n_{b},n_{c}\in\mathbb{R}$, and suppose that the quadratic equation
$z^{2}+(1+n_{b})z-n_{c}=0$ (3.10)
has two distinct real roots $\alpha<\beta$. Then under Assumption 3.4, the assertions in (i) hold with $\lambda_{0}=0$ and $\alpha p<\theta<\beta p$, where the dependencies of $N$ are replaced by $d,q,p,\theta,n_{b},n_{c},\nu,K$, and $K_{0}$.
@@
Here, in the case when Assumption 3.4 holds, the ratio condition (3.8) is still satisfied. This implies that Assumption $2.2$ $(\rho_{0},\gamma_{0})$ or $2.3$ $(\rho_{0},\gamma_{0})$ in [15] is satisfied. Hence, by [15, Theorem 2.6], there exists $\lambda_{0}\geq 0$ such that for any $\lambda\geq\lambda_{0}$, one can find a solution $u\in\mathbb{H}_{p,\theta,\omega}^{1}(T)$ to (3.11) satisfying
@@
where $\Lambda\geq 0$, and $g$ is defined as (3). Here, $h_{r}(t,x):=h(t,x/r)$ for any function $h$ on $\Omega_{T}$. In particular, $v\in\mathcal{H}_{p,\omega}^{1}(T)$ where $\mathcal{H}_{p,\omega}^{1}(T)$ is the set of functions such that $v,D_{x}v\in L_{p}((-\infty,T)\times\mathbb{R}^{d},\omega(t)dtdx)$ and $a_{0,r}h_{t}\in L_{p}((-\infty,T),\omega dt;H_{p}^{-1}(\mathbb{R}^{d}))$. Since $g\in L_{p}((-\infty,T)\times\mathbb{R}^{d},\omega(t)dtdx)$, by [8, Theorem 6.2], we have a solution to (3) in the space $W_{p,\omega}^{1,2}(T)$, say $\tilde{v}$. Since $\tilde{v}$ is also in the space $\mathcal{H}_{p,\omega}^{1}(T)$, the uniqueness result in $\mathcal{H}_{p,\omega}^{1}(T)$ (see also [15, Remark 3.7]) yields that $v=\tilde{v}\in W_{p,\omega}^{1,2}(T)$, and
@@
Before considering the general case $q\neq p$, we remark that $u$ is independent of $p,\theta$, and $\omega$ when $f\in C_{c}^{\infty}(\Omega_{T})$. In other words, for $p_{i}\in(1,\infty)$, $\omega_{i}\in A_{p_{i}}(\mathbb{R})$, and $\theta_{i}\in\mathbb{R}$ ($\theta_{i}\in(\alpha p_{i},\beta p_{i})$ for the case $(ii)$), if $u_{i}\in\mathfrak{H}_{p_{i},\theta_{i},\omega_{i}}^{2}(T)$ are two solutions to (1.1) with $f\in C_{c}^{\infty}(\Omega_{T})$, then $u_{1}=u_{2}$. This fact easily follows from [15, Remark 4.6], where the independence for divergence form equation (3.11) is addressed.
Next, we deal with the existence when $q\neq p$. Let $f\in C_{c}^{\infty}(\Omega_{T})$, $\omega^{\prime}\in A_{p}(\mathbb{R})$, and take a solution $u\in\mathfrak{H}_{p,\theta,\omega^{\prime}}^{2}(T)$. Since $u$ is independent of $\omega^{\prime}$, it satisfies the estimate (3.5) with $q=p$ and any $\omega^{\prime}\in A_{p}(\mathbb{R})$. This and the extrapolation theorem (see e.g. [8, Theorem 2.5]) yield that $u\in\mathfrak{H}_{q,p,\theta,\omega}^{2}(T)$, and (3.5) for general $q\neq p$ and $\omega\in A_{q}(\mathbb{R})$.
Lastly, we consider the uniqueness result. Let $u\in\mathfrak{H}_{q,p,\theta,\omega}^{2}(T)$ be a solution to (1.1) with $f=0$. As in the first case $p=q$, we may assume that $a_{dd}=1$, and $u$ satisfies the divergence form equation (3.11). By the uniqueness result from [15, Theorem 2.6], we have $u=0$. The lemma is proved. ∎
@@
Let $\varepsilon_{0}\in(0,1)$. We consider a partition of unity with weight. By [24, Lemma 5.6], there exists $\rho_{0}=\rho_{0}(\varepsilon_{0})\in(1/2,1)$, and nonnegative $\eta_{k}\in C_{c}^{\infty}(\mathbb{R}^{d}_{+})$ such that
@@
To prove Theorem 2.6, it suffices to repeat the proof of [15, Theorem 2.9], where we adopt the idea of the proof of [29, Theorem 2.1]. For the reader’s convenience, we provide the full proof.
@@ 4. Elliptic equations
Now we apply [5, Theorem 2.5] to obtain that
@@
for some $n_{b},n_{c}\in\mathbb{R}$. Assume that the quadratic equation
$z^{2}+(1+n_{b})z-n_{c}=0$ (4.6)
has two distinct real roots $\alpha<\beta$. Then for any $\theta\in\mathbb{R}\setminus[\alpha p,\beta p]$ and $f\in L_{p,\theta}$, there is a unique solution $u\in H_{p,\theta}^{2}$ to
@@
By [15, Theorem 2.14], there is $\lambda_{0}\geq 0$ such that for $\lambda\geq\lambda_{0}$, one can find a solution $u\in H_{p,\theta}^{1}$ to (4.11). Moreover, for this solution, we have
@@
where $\Lambda\geq 0$ and $g$ is given by (4.3). Since $g\in L_{p}(\mathbb{R}^{d})$, one can apply [5, Theorem 2.5] to obtain a solution $\tilde{v}$ to (4.13) in $W_{p}^{2}(\mathbb{R}^{d})$. Since (4.13) can be understood in divergence form, the uniqueness result in $W_{p}^{1}(\mathbb{R}^{d})$ leads to $v=\tilde{v}\in W_{p}^{2}(\mathbb{R}^{d})$, and
@@
5. Optimality of the range of weights
@@
1. Elliptic equations when $d=1$.
@@
2. Parabolic equations when $d=1$.
@@
3. Parabolic and elliptic equations when $d\geq 2$.
In this step, we prove that when $d\geq 2$ and $\lambda=0$, the range $(\alpha p,\beta p)$ is sharp in Theorems 2.4 and 2.9. As in 2, by investigating the proof of Theorem 2.9, it suffices to show that the range is sharp for the elliptic equations.
@@
where $p^{\prime}=p/(p-1)$. One can also notice that $v(x):=\zeta(x_{d})u(x/r)$ satisfies (4.2) with $f_{r}=0$, where $\zeta\in C_{c}^{\infty}((2,3))$ is a standard nonnegative cutoff function. Then by the uniqueness result of [5, Theorem 2.5], we have $v\in W_{p}^{2}(\mathbb{R}^{d})$. Thus, we can repeat the proof of Lemma 4.1 to obtain that
@@
Since $2\theta/p>2\beta=2\sqrt{c}$, by (5),
@@
Thus, by (5), (5), and (5.9), we have $u\zeta_{n}\in H_{p,\theta}^{2}$.
@@
where $B_{n}:=\mathbb{R}^{d-1}\times\left((1/(2n),1/n)\cup(n,2n)\right)$, and $N$ is independent of $n$. Combining (5), (5.9), and (5),

### v2

@@ Abstract.
Corresponding author. J. Ryu was supported by a KIAS Individual Grant (MG101501) at Korea Institute for Advanced Study and by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2026-25477288).
@@ 1. Introduction
Observe that equations (1.1) and (1.2) are stated without boundary conditions, as their solvability in weighted Sobolev spaces does not require such conditions. Nevertheless, it is possible to study these equations under certain boundary conditions. In such cases, additional assumptions on the forcing term $f$ or on the structure of the equations are necessary (see e.g. [43, Remark 4.7]). In this paper, however, our goal is to address the equations under minimal assumptions.
@@
where $\alpha>0$, and $H:\Omega_{T}\times\mathbb{R}^{d}\to\mathbb{R}$ is a given smooth Hamiltonian. Here, equation (1.1) corresponds to the special case $H=0$ and $\alpha=2$. Hamilton–Jacobi equations play an important role in optimal control, differential games, and related areas. In particular, regularity theory is a fundamental subject in nonlinear PDEs. We refer the reader to [2, 16] for more information. Furthermore, the equations are derived from a variety of problems. For instance, (1.2) is obtained in the linearization of the nonlinear and degenerate Loewner–Nirenberg problem
@@
The paper is organized as follows. In Section 2, we introduce the weighted Sobolev spaces, assumptions, and our main results. In Section 3, we provide the proofs of the main results on the parabolic problem (Theorems 2.4 and 2.7). In Section 4, we prove Theorem 2.10, which concerns the elliptic problem. We address the optimality of our main results in Section 5. Lastly, in Section 6, we show that our main results can be applied to degenerate viscous Hamilton–Jacobi equations when the Hamiltonian is of certain specific forms.
@@ 2.1. Notation and function spaces
We remark that $H_{p,\theta}^{n}$ corresponds to $H_{p,\theta+d-1}^{n}$ in [30]. Moreover, from [30, Corollary 2.3], the norm (2.1) is equivalent to
@@
To the best of our knowledge, the spaces $H_{p,\theta}^{n}$ were firstly introduced in [39, Section 2.6.3] for $p=2$ and $\theta=1$. They were generalized in a unified manner for $p\in(1,\infty)$ and $\theta\in\mathbb{R}$ in [30] to study stochastic partial differential equations (SPDEs). We refer the reader to [29, 33, 34]. See also [25, 28, 9, 42, 38], where second-order nondegenerate equations are studied.
@@
By [30, Theorem 1.19, Remark 5.5], the spaces $C_{c}^{\infty}(\mathbb{R}^{d}_{+})$, $C_{c}^{\infty}((-\infty,T)\times\mathbb{R}^{d}_{+})$, and $C_{c}^{\infty}((-\infty,T]\times\mathbb{R}^{d}_{+})$ are dense in $H_{p,\theta}^{n}$, $\mathbb{H}_{q,p,\theta,\omega}^{n}(T)$, and $\mathfrak{H}_{q,p,\theta,\omega}^{2}(T)$, respectively. For the Cauchy problems, the space $\mathring{\mathfrak{H}}_{q,p,\theta,\omega}^{2}(0,T)$ is defined by the closure of the set of functions $u\in C_{c}^{\infty}([0,T]\times\mathbb{R}^{d}_{+})$ with $u(0,\cdot)=0$, equipped with the norm
@@ 2.2. Parabolic equations
We provide some comments on the components in the statement of Theorem 2.4.
$(i)$ The parameter $\rho_{0}$ is closely related to the construction of a partition of unity in the weighted spaces. Unlike the unweighted case, weighted derivatives of the partition of unity can be made arbitrarily small by choosing $\rho_{0}$ sufficiently close to $1$ (see (3.18)). This smallness is essential for controlling the localization errors in the a priori estimates.
$(ii)$ The smallness of $\gamma_{0}$ implies that the coefficients have small bounded mean oscillations in the relevant spatial variables.
$(iii)$ The range of $\theta$ in $(ii)$ of Theorem 2.4 is determined by $n_{b}$ and $n_{c}$. Roughly, as in the ODE case (1.6), the behavior of the solution is influenced by $a_{dd},b_{d}$, and $c$. We also note that (2.7) is equivalent to (1.7) when $d=1$ and the coefficients are constant.
$(iv)$ In $(i)$ of the theorem, the lower order coefficients $(b_{i})$ and $c$ are assumed to be merely bounded measurable, and $\theta\in\mathbb{R}$ is arbitrary, although we need to choose a sufficiently large $\lambda_{0}\geq 0$. On the other hand, in $(ii)$, we can take $\lambda=0$ by imposing ratio conditions on the lower-order coefficients and restricting the range of $\theta$.
Remark 2.6.
In [17, Theorem 2.6 $(ii)$], the corresponding divergence form equation was studied under the assumption that either
@@ 2.3. Elliptic equations
$(ii)$ Let $n_{b},n_{c}\in\mathbb{R}$. Suppose that the quadratic equation (2.7) has two distinct real roots $\alpha<\beta$. Then there exist
@@
Remark 2.11.
By dividing (1.2) by $x_{d}^{2}$, the equation can be viewed as a nondegenerate equation with singular lower-order coefficients satisfying a growth condition. Such equations were studied in [9, Theorem 2.3], which corresponds to a special case of Theorem 2.10 with $n_{b}=n_{c}=0$,
@@ 3. Parabolic equations
In this section, we deal with parabolic equations and give the proofs of Theorems 2.4 and 2.7.
@@
Using (2.2), one can show that there is $R_{0}>0$ such that the (unweighted) oscillation of $x_{d}^{2}a_{ij,r}$ on $Q_{R}(t,x):=(t-R^{2},t)\times B_{R}(x)$ is less than $N\gamma_{0}$ if $x_{d}\in(1,4)$ and $R\leq R_{0}$. By [10, Theorem 6.3], there is $\Lambda_{0}=\Lambda_{0}(d,p,\nu,K_{0})\geq 0$ such that if $\Lambda=\Lambda_{0}$, then we can apply the $W_{p,\omega}^{1,2}$-estimate for the equation (3.2) to get
@@
where $L_{p,\omega}(T):=L_{p}((-\infty,T)\times\mathbb{R}^{d},\omega(t)dtdx)$. Here, we remark that the arguments in [10] still work for equations with general coefficients $a_{0}$ and $c_{0}$, and thus we can still utilize the estimate established in the paper (see also [17, Remark 3.7]). Note that for any $\gamma\in\mathbb{R}$,
@@
(ii) Let $n_{b},n_{c}\in\mathbb{R}$, and suppose that the quadratic equation (2.7) has two distinct real roots $\alpha<\beta$. Then under Assumption 3.4, the assertions in (i) hold with $\lambda_{0}=0$ and $\alpha p<\theta<\beta p$, where the dependencies of NN are replaced by $d,q,p,\theta,n_{b},n_{c},\nu,K$, and $K_{0}$.
@@
Here, in the case when Assumption 3.4 holds, the ratio condition (3.8) is still satisfied. This implies that Assumption $2.2$ $(\rho_{0},\gamma_{0})$ or $2.3$ $(\rho_{0},\gamma_{0})$ in [17] is satisfied. Hence, by [17, Theorem 2.6], there exists $\lambda_{0}\geq 0$ such that for any $\lambda\geq\lambda_{0}$, one can find a solution $u\in\mathbb{H}_{p,\theta,\omega}^{1}(T)$ to (3.10) satisfying
@@
where $\Lambda\geq 0$, and $g$ is defined as (3). Here, $h_{r}(t,x):=h(t,x/r)$ for any function $h$ on $\Omega_{T}$. In particular, $v\in\mathcal{H}_{p,\omega}^{1}(T)$ where $\mathcal{H}_{p,\omega}^{1}(T)$ is the set of functions such that $v,D_{x}v\in L_{p}((-\infty,T)\times\mathbb{R}^{d},\omega(t)dtdx)$ and $a_{0,r}h_{t}\in L_{p}((-\infty,T),\omega dt;H_{p}^{-1}(\mathbb{R}^{d}))$. Since $g\in L_{p}((-\infty,T)\times\mathbb{R}^{d},\omega(t)dtdx)$, by [10, Theorem 6.2], we have a solution to (3) in the space $W_{p,\omega}^{1,2}(T)$, say $\tilde{v}$. Since $\tilde{v}$ is also in the space $\mathcal{H}_{p,\omega}^{1}(T)$, the uniqueness result in $\mathcal{H}_{p,\omega}^{1}(T)$ (see also [17, Remark 3.7]) yields that $v=\tilde{v}\in W_{p,\omega}^{1,2}(T)$, and
@@
Before considering the general case $q\neq p$, we remark that $u$ is independent of $p,\theta$, and $\omega$ when $f\in C_{c}^{\infty}(\Omega_{T})$. In other words, for $p_{i}\in(1,\infty)$, $\omega_{i}\in A_{p_{i}}(\mathbb{R})$, and $\theta_{i}\in\mathbb{R}$ ($\theta_{i}\in(\alpha p_{i},\beta p_{i})$ for the case $(ii)$), if $u_{i}\in\mathfrak{H}_{p_{i},\theta_{i},\omega_{i}}^{2}(T)$ are two solutions to (1.1) with $f\in C_{c}^{\infty}(\Omega_{T})$, then $u_{1}=u_{2}$. This fact easily follows from [17, Remark 4.6], where the independence for divergence form equation (3.10) is addressed.
Next, we deal with the existence when $q\neq p$. Let $f\in C_{c}^{\infty}(\Omega_{T})$, $\omega^{\prime}\in A_{p}(\mathbb{R})$, and take a solution $u\in\mathfrak{H}_{p,\theta,\omega^{\prime}}^{2}(T)$. Since $u$ is independent of $\omega^{\prime}$, it satisfies the estimate (3.5) with $q=p$ and any $\omega^{\prime}\in A_{p}(\mathbb{R})$. This and the extrapolation theorem (see e.g. [10, Theorem 2.5]) yield that $u\in\mathfrak{H}_{q,p,\theta,\omega}^{2}(T)$, and (3.5) for general $q\neq p$ and $\omega\in A_{q}(\mathbb{R})$.
Lastly, we consider the uniqueness result. Let $u\in\mathfrak{H}_{q,p,\theta,\omega}^{2}(T)$ be a solution to (1.1) with $f=0$. As in the first case $p=q$, we may assume that $a_{dd}=1$, and $u$ satisfies the divergence form equation (3.10). By the uniqueness result from [17, Theorem 2.6], we have $u=0$. The lemma is proved. ∎
@@
Let $\varepsilon_{0}\in(0,1)$. We consider a partition of unity with weight. By [26, Lemma 5.6], there exists $\rho_{0}=\rho_{0}(\varepsilon_{0})\in(1/2,1)$, and nonnegative $\eta_{k}\in C_{c}^{\infty}(\mathbb{R}^{d}_{+})$ such that
@@
To prove Theorem 2.7, it suffices to repeat the proof of [17, Theorem 2.9], where we adopt the idea of the proof of [31, Theorem 2.1]. For the reader’s convenience, we provide the full proof.
@@ 4. Elliptic equations
Now we apply [7, Theorem 2.5] to obtain that
@@
for some $n_{b},n_{c}\in\mathbb{R}$. Assume that the quadratic equation (2.7) has two distinct real roots $\alpha<\beta$. Then for any $\theta\in\mathbb{R}\setminus[\alpha p,\beta p]$ and $f\in L_{p,\theta}$, there is a unique solution $u\in H_{p,\theta}^{2}$ to
@@
By [17, Theorem 2.14], there is $\lambda_{0}\geq 0$ such that for $\lambda\geq\lambda_{0}$, one can find a solution $u\in H_{p,\theta}^{1}$ to (4.10). Moreover, for this solution, we have
@@
where $\Lambda\geq 0$ and $g$ is given by (4.3). Since $g\in L_{p}(\mathbb{R}^{d})$, one can apply [7, Theorem 2.5] to obtain a solution $\tilde{v}$ to (4.12) in $W_{p}^{2}(\mathbb{R}^{d})$. Since (4.12) can be understood in divergence form, the uniqueness result in $W_{p}^{1}(\mathbb{R}^{d})$ leads to $v=\tilde{v}\in W_{p}^{2}(\mathbb{R}^{d})$, and
@@
5. Optimality of the weights when $\lambda=0$
@@
5.1. Elliptic equations when $d=1$
@@
5.2. Parabolic equations when $d=1$
@@
5.3. Parabolic and elliptic equations when $d\geq 2$
In this step, we prove that when $d\geq 2$ and $\lambda=0$, the range $(\alpha p,\beta p)$ is sharp in Theorems 2.4 and 2.10. As in 2, by investigating the proof of Theorem 2.10, it suffices to show that the range is sharp for the elliptic equations.
@@ 5.3. Parabolic and elliptic equations when $d\geq 2$
where $p^{\prime}=p/(p-1)$. One can also notice that $v(x):=\zeta(x_{d})u(x/r)$ satisfies (4.2) with $f_{r}=0$, where $\zeta\in C_{c}^{\infty}((2,3))$ is a standard nonnegative cutoff function. Then by the uniqueness result of [7, Theorem 2.5], we have $v\in W_{p}^{2}(\mathbb{R}^{d})$. Thus, we can repeat the proof of Lemma 4.1 to obtain that
@@
Since $2\theta/p>2\beta=2\sqrt{c}$, by (5.3),
@@
Thus, by (5.3), (5.3), and (5.9), we have $u\zeta_{n}\in H_{p,\theta}^{2}$.
@@
where $B_{n}:=\mathbb{R}^{d-1}\times\left((1/(2n),1/n)\cup(n,2n)\right)$, and $N$ is independent of $n$. Combining (5.3), (5.9), and (5.3),